# Resilience of Large Language Models for Noisy Instructions

### v1

@@ 2 Related Work
The progression of general-purpose Large Language Models (LLMs) such as ChatGPT Achiam et al. (2023), Gemini Team et al. (2023), LLaMa Touvron et al. (2023), Mistral Jiang et al. (2023a), and Gemma Mesnard et al. (2024) has facilitated a myriad of real-world applications. These advancements are attributed to their capabilities in managing long-range textual dependencies, enhancing contextual comprehension, and displaying a remarkable ability to adapt to a wide array of tasks with minimal need for detailed, task-specific training. Meantime, several recent studies have demonstrated that the user prompt significantly influences task performance, highlighting its indispensable role in the process Wang et al. (2023); Zhu et al. (2023). The following will introduce studies on prompt sensitivity and noisy text reconstruction as related work.
@@ 2.1 Instruction Sensitivity
Pre-trained large language models exhibit performance variability even with semantically similar inputs. SeaEval Wang et al. (2023) demonstrates that across five different input templates, performance can fluctuate between 5% and 10%, depending on the dataset. Similar observations are reported in Sclar et al. (2023), highlighting this instability. Moreover, introducing brief sentences such as "Let’s think step by step" can significantly enhance performance on reasoning tasks, further underscoring the LLMs’ sensitivity Kojima et al. (2022). Leveraging this characteristic, various studies concentrate on decomposing and crafting improved prompts to efficiently tackle tasks. Zhou et al. (2023) proposes an additional LLM as a prompt engineer to automatically create prompts that enhance performance. Alternatively, some approaches advocate for the use of search Prasad et al. (2023) or optimization techniques Khot et al. (2022); Hao et al. (2022); Prasad et al. (2023) to identify superior instructions, replacing those that are less effective.
@@ 3.1.1 Method
Using the provided dataset of clean and predicted transcripts, we divide them into four distinct categories based on their Word-Error-Rate (WER): less than 10%, 10%-20%, 20%-30%, and 30%-40%, allocating 80,000 samples to each category. By leveraging this paired data, we finetune Tiny-Llama-1.1B-Chat Zhang et al. (2024) models for each category to learn the underlying patterns of ASR errors. We use the trained error generation model to introduce ASR errors into each question sentence from the MMLU dataset independently. This approach produces a varied WER (Word Error Rate), as illustrated in Figure 3. This noise injection method is consistently applied across all types of noise, except for the destructive content. The same set of 1,000 questions is selected for five noisy instructions to benchmark the performance.
@@ 3.1.2 Discussion
Although LLMs have numerous applications in processing spoken content, they lack robustness against errors introduced by ASR systems. Consequently, there is a critical need for the development of LLMs that are resilient to ASR errors, as well as the creation of comprehensive speech-to-text foundation models that can directly handle speech inputs Chu et al. (2023); Tang et al. (2024).
@@ 4.3 Discussion on Efficacy
In this section, we explore how effective Large Language Models (LLMs) are at mitigating the impact of noisy instructions. However, there are two major drawbacks: 1) open-sourced models generally perform poorly in this task and 2) there is an extra computational cost associated with processing requests. Therefore, there is a need for a lightweight model that is task-agnostic for noisy instruction correction. During our research, we explored fine-tuning an LLM with 1.1B model Zhang et al. (2024) sizes using synthesized data to recover clean instructions from noisy ones. However, we found it is challenging for the model to grasp the real intent behind instructions and as a result, unable to performance error corrections accurately. This difficulty is attributed to the constraints imposed by the model size. Therefore, efficient correction of instructions require further investigation, which holds significant use cases like defending adversarial attacks and system integration (e.g. ASR and OCR).
@@ 5 Conclusion
Second, our analysis and error types are limited to English benchmarks without extension to multilingual scenarios. The problem becomes more complex as each language has its own uniqueness. Moreover, coding-switching noise introduce further complexities. The assessment of LLM’s resilience to noisy instructions in multilingual scenarios is an area needs future explorations.
Last, in this study, we focus on five types of noise rooted from system integration (ASR, OCR) and user interactions (Typographical, Grammatical, and Distraction Content). While comprehensive within its defined scope, our work does not encompass all possible sources of noise that could affect LLM performance. For instance, semantic ambiguities or stylistic variations, which could significantly impact the interpretation and processing capabilities of LLMs, are not investigated in details.

### v2

@@ 2 Related Work
The progression of general-purpose Large Language Models (LLMs) such as ChatGPT Achiam et al. (2023), Gemini Team et al. (2023), LLaMa Touvron et al. (2023), Mistral Jiang et al. (2023a), and Gemma Mesnard et al. (2024) has facilitated a myriad of real-world applications. These advancements are attributed to their capabilities in managing long-range textual dependencies, enhancing contextual comprehension, and displaying a remarkable ability to adapt to a wide array of tasks with minimal need for detailed, task-specific training. Meantime, several recent studies have demonstrated that the user prompt significantly influences task performance, highlighting its indispensable role in the process Wang et al. (2024b); Zhu et al. (2023). The following will introduce studies on prompt sensitivity and noisy text reconstruction as related work.
@@ 2.1 Instruction Sensitivity
Pre-trained large language models exhibit performance variability even with semantically similar inputs. SeaEval Wang et al. (2024b) demonstrates that across five different input templates, performance can fluctuate between 5% and 10%, depending on the dataset. Similar observations are reported in Sclar et al. (2023), highlighting this instability. Moreover, introducing brief sentences such as "Let’s think step by step" can significantly enhance performance on reasoning tasks, further underscoring the LLMs’ sensitivity Kojima et al. (2022). Leveraging this characteristic, various studies concentrate on decomposing and crafting improved prompts to efficiently tackle tasks. Zhou et al. (2023) proposes an additional LLM as a prompt engineer to automatically create prompts that enhance performance. Alternatively, some approaches advocate for the use of search Prasad et al. (2023) or optimization techniques Khot et al. (2022); Hao et al. (2022); Prasad et al. (2023) to identify superior instructions, replacing those that are less effective.
@@ 3.1.1 Method
Using the provided dataset of clean and predicted transcripts, we divide them into four distinct categories based on their Word-Error-Rate (WER): less than 10%, 10%-20%, 20%-30%, and 30%-40%, allocating 80,000 samples to each category. By leveraging this paired data, we finetune Tiny-Llama-1.1B-Chat Zhang et al. (2024a) models for each category to learn the underlying patterns of ASR errors. We use the trained error generation model to introduce ASR errors into each question sentence from the MMLU dataset independently. This approach produces a varied WER (Word Error Rate), as illustrated in Figure 3. This noise injection method is consistently applied across all types of noise, except for the destructive content. The same set of 1,000 questions is selected for five noisy instructions to benchmark the performance.
@@ 3.1.2 Discussion
Although LLMs have numerous applications in processing spoken content, they lack robustness against errors introduced by ASR systems. Consequently, there is a critical need for the development of LLMs that are resilient to ASR errors, as well as the creation of comprehensive speech-to-text foundation models that can directly handle speech inputs Chu et al. (2023); Tang et al. (2024); Wang et al. (2024c); Zhang et al. (2024b).
@@ 4.3 Discussion on Efficacy
In this section, we explore how effective Large Language Models (LLMs) are at mitigating the impact of noisy instructions. However, there are two major drawbacks: 1) open-sourced models generally perform poorly in this task and 2) there is an extra computational cost associated with processing requests. Therefore, there is a need for a lightweight model that is task-agnostic for noisy instruction correction. During our research, we explored fine-tuning an LLM with 1.1B model Zhang et al. (2024a) sizes using synthesized data to recover clean instructions from noisy ones. However, we found it is challenging for the model to grasp the real intent behind instructions and as a result, unable to performance error corrections accurately. This difficulty is attributed to the constraints imposed by the model size. Therefore, efficient correction of instructions require further investigation, which holds significant use cases like defending adversarial attacks and system integration (e.g. ASR and OCR).
@@ 5 Conclusion
6 Acknowledgement
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore.
@@ 6 Acknowledgement
Second, our analysis and error types are limited to English benchmarks without extension to multilingual scenarios. The problem becomes more complex as each language has its own uniqueness. Moreover, coding-switching noise introduces further complexities. The assessment of LLM’s resilience to noisy instructions in multilingual scenarios is an area needs future explorations.
Last, in this study, we focus on five types of noise rooted from system integration (ASR, OCR) and user interactions (Typographical, Grammatical, and Distraction Content). While comprehensive within its defined scope, our work does not encompass all possible sources of noise that could affect LLM performance. For instance, semantic ambiguities, stylistic variations or multilingualism, which could significantly impact the interpretation and processing capabilities of LLMs, are not investigated in detail Lovenia et al. (2024); Wang et al. (2024a); Lin et al. (2024).